# DPOD: Domain-Specific Prompt Tuning for Multimodal Fake News Detection

### v3

@@ 4.2 Stage 2: Computing Semantic Domain Vectors
In this stage, we create semantic domain vectors which represent how similar each domain is to all the domains in the training dataset. If the number of domains in the training dataset $\mathcal{S}_{train}$ is $n$, we create $n$ semantic domain vectors each having $n$ indices. Let us take the “politics” domain in the training dataset as an example. To compute the semantic domain vector for “politics”, we pass all its images and corresponding text captions, i.e., $\{({I}_{k},{T}_{k})\},k=1,2,...,{n}_{{D}_{i}}$ through the frozen encoders of the A-CLIP trained in Stage 1 to get $\{({\hat{I}}_{k},{\hat{T}}_{k})\}$. Here, ${n}_{{D}_{i}}$ denotes the number of training samples in the politics domain. We take a Hadamard product of the image and text embedding $\{({\hat{I}}_{k},{\hat{T}}_{k})\}$ to get ${n}_{{D}_{i}}$ joint embedding. Then, we compute the mean across all the sample joint embedding to get the final mean joint embedding $\mathcal{J}_{{D}_{i}}$, particular to a domain ${D}_{i}$ (here, “politics”) as shown below:
@@
This is illustrated in Fig. 1(a) Finally, to create the semantic domain vector ${w}_{{D}_{i}}$, we take the cosine similarity between this joint embedding $\mathcal{J}_{{D}_{i}}$ and those of the other training domains $\mathcal{J}_{{D}_{k}},k=1,2,...,n$. Hence, the $j^{th}$ index of the ${w}_{{D}_{i}}$ vector will be the cosine similarity of $\mathcal{J}_{{D}_{i}}$ with $\mathcal{J}_{{D}_{j}}$, ($j\in k$) as shown in eq. (6).
@@ 4.3 Stage 3: Domain-Specific Prompt-tuning with Out-of-Domain data
In this stage, we utilize the semantic domain vectors to learn domain-specific prompt embedding that help to incorporate domain-specific information in the model while classifying a news item as true or fake. We initialize a prompt prefix with the standard text “A photo of” [18]. However, we make this prefix learnable (similar to [28]), since it may not be the best choice for our application. We denote these generic learnable prompts as $\{{V}_{1},{V}_{2},{V}_{3}\}$. To incorporate the domain information, we pass the corresponding semantic domain vector ${w}_{{D}_{i}}$ (obtained in the previous stage) through a linear layer, denoted by $\mathcal{F}_{\theta}$. This layer $\mathcal{F}_{\theta}$ projects the ${n}_{{D}_{i}}$-dimensional domain vector to a 512-dimensional domain prompt token $\mathcal{V}_{{D}_{i}}$, which is then appended to the prompts $\{{V}_{1},{V}_{2},{V}_{3}\}$. This forms the whole learnable prompt structure, where $\{{V}_{1},{V}_{2},{V}_{3}\}$ is trained in a domain-agnostic manner, whereas, $\mathcal{V}_{{D}_{i}}$ is trained by conditioning on the domain of the particular training sample. This entire prompt is then prepended to the text caption ${T}$ and passed through the frozen A-CLIP text encoder to get a text embedding $\mathcal{\hat{T}}$ as shown in eq. (7).
@@
The Classifier Network consists of two fully connected layers which project $\mathcal{\hat{J}}$ to $\mathcal{\hat{X}}$. Inspired by [7], we also introduce a residual connection from $\mathcal{\hat{J}}$ to $\mathcal{\hat{X}}$, the output of which finally passes through two fully connected layers to give a final prediction $\hat{y}$. The whole network is trained in an end-to-end manner using a Binary Cross-Entropy loss as shown in eq. (8).
@@
Difference with other prompt-learning techniques. As discussed in the Related Work section, there has been a recent few works on domain-based prompt tuning [8, 21]. But there are few important differences between these techniques and the proposed one:
1) In the existing works, “domain” or “style” refers to changes in image features. Contrary to this, in our case, domains (such as “politics”, “law-crime”, etc.) do not have significant change in image features (since, images from multiple domains will contain similar objects like humans, etc.), but are subjective to human interpretation of both the images and the texts jointly.
@@
Table 1: Comparison of the performances of our proposed method against existing approaches on the NewsCLIPpings dataset where we train and test on all the domains.
Methods Backbone Accuracy (%) DT-Transformer [17] CLIP ViT-B/32 65.67 MNSL [25] CLIP 68.20 DCD [10] CLIP 65.40 NewsCLIPpings [15] CLIP ViT-B/32 60.23 CLIP-FT [1] CLIP ViT-B/16 64.79 CLIP ViT-B/32 65.10 SSDL [16] CLIP ViT-B/16 65.00 CLIP RN-50 71.00 DPOD (Ours) CLIP ViT-B/16 70.49 DPOD (Ours) CLIP ViT-B/32 70.81 CLIP RN-50 74.43
@@ 5 Experimental Evaluation
We evaluate the effectiveness of the proposed framework on the standard large-scale benchmark, namely NewsCLIPpings dataset [15], which is created from the Visual News dataset [14]. It contains news articles from sources like: The Guardian, BBC, USA Today, and The Washington Post. It has a total of $71,072$ train, $7,024$ validation and $7,264$ test examples. It contains news articles belonging to a total of 54 domains with varied amounts of training data across each of the domains for example, politics (1396 samples), law-crime (4348 samples), sports (2786 samples) etc., effectively relecting the imbalance problem.
Implementation Details. CLIP “ViT-B/32”[18] is the backbone of all our experiments. We use accuracy as the evaluation metric as in [1, 16], as the data has equal distribution of true and fake news. Our codes are implemented in PyTorch and uses a single NVIDIA GeForce RTX 3080 Ti GPU. ADAM optimizer with weight decay and stability constant is used for the training. All the results are reported using the same set of parameters, e.g., $\beta=1.5$, $lr=1e-4$ and $batchsize=64$ across all dataset splits. The values of hyperparameters are obtained during the validation process against the constant validation dataset containing all-domain data. These values were then kept fixed for the rest of the experiments. Stage 1 is run for 40 epochs, which remains fixed even if the desired domain changes. We initialize our domain-agnostic prompts $\{V_{1},V_{2},V_{3}\}$ required in Stage 3, as “A photo of” which is then learnt through training. After obtaining the output probability at the end of Stage 3, we use Sigmoid function to convert the output into a probability score $\hat{y}$ between 0 to 1. We train Stage 3 for 30 epochs. At the time of inference, the data having output greater than or equal to $0.5$ is assigned the label 1, i.e., Fake and the others are assigned label 0, i.e., True. Now, we describe the experiments conducted to evaluate the effectiveness of the proposed DPOD framework. We would specifically like to answer the following research questions:
  1. 1.
    How does the model perform for the standard setting, where all the domain data is used for both training and testing?
  2. 2.
    How well does the proposed model perform for a desired target domain under the imbalance scenario?
  3. 3.
    Can our model generalize to completely unseen target domains?
  4. 4.
    Can the model handle inconsistencies in domain labels and how well does it understand about the domains?
  5. 5.
    Are all the components in the model useful?
@@
Radio(470)
Education(376) Money(224) Travel(526)
@@
Evaluation on standard setting for all domains. Our proposed DPOD is a generalized framework, where we train our model with data from all the diverse domains available. Along with evaluating on all the domains combined, the framework can also generalize to domains of interest, by leveraging information of other domains. First, we evaluate our model on image-text pairs available from all the domains combined, and report the results in Table 1. Here, we assume that we know the domain label of the test data. The experiment is run $5$ times with random sampling of training data, and the mean performance is reported. We compare our proposed framework with the previous SOTA approaches on NewsCLIPpings dataset, namely, CLIP-FT [1], SSDL [16], DT-Transformer [17], MNSL[25], DCD[10], NewsCLIPpings[15]. For CLIP-FT, the entire CLIP model is fine-tuned along with two additional layers with the available training data. This is the performance without using the open-domain data. For all the other approaches, we directly report the accuracies reported in their papers. For the scenario where domain knowledge of the test sample is unknown, our model gives an accuracy of $69.13\%$, which is comparable to the performance achieved when the test domain information is known.
For fair comparison, we also perform our experiments with different variants of the CLIP model, namely, ViT-B/32, ViT-B/16 and RN-50. We observe that the proposed framework consistently outperforms the SOTA by a significant margin when similar backbones are used for comparison, in both the settings.
@@
Figure 2: (a) Computation of the joint embedding of the images and text for a particular domain. These embedding are used in Stage 2 and 3 of our DPOD framework. (b) Similarity of the learnt domain-specific prompts of different domains. We observe that the sets Football, Sport, Sports and Healthcare Network, Healthcare Medicine have similar domain-specific prompts among themselves, but they have low similarity with each other and also with Music.
Evaluation on a specific domain. Now, we consider the case where an individual might be interested in news from a specific domain. We report in Table 2 how well our model utilizes out-of-domain samples in order to predict the veracity of the news belonging to a desired domain. To highlight the problem of data imbalance across different domains, we consider $5$ domains with the most datapoints, namely, Sport, Law-crime, International Relations, World, Arts&Culture, along with $5$ domains with the least data samples, namely, Fashion. TV-Radio, Education, Money, Travel. In the absence of comparative approaches, we experiment with the following strong baselines:
1) Clip-FT [1]: without the open-domain information for fair comparison with the other approaches.
@@
2) CoOp [28]: First, the prompts $\{V_{1},V_{2},V_{3}\}$ are learnt using training data from all the domains. Then, these prompts are frozen for evaluation on the target domains. For fair comparison with our approach, CoOp is also used with A-CLIP. All the models here are trained on the entire dataset. From Table 2, we observe that DPOD consistently outperforms the other compared methods for all the individual domains, irrespective of the number of training examples available for the particular domain, highlighting the effectiveness of the proposed domain-specific prompt learning framework.
@@
Table 3: Comparison of accuracies (in %) where we train the model on 90% domains and evaluate its performance on the rest 10% unseen domains.
Methods Politics Law Crime Arts & Cul. Sports Music
CLIP-FT 68.31 61.65 64.87 71.05 63.33
DPOD 72.29 64.63 71.48 73.18 66.67
@@
Generalization to unseen domains. Till now, we have assumed that the test domain is already present during training. Here, we evaluate the performance of our model when presented with completely new domains not included in the training set. For this, we train our model using samples from 90% of the domains, and test it on the remaining 10% unseen domains, using the most similar domain-specific prompt (based on cosine similarity). For a fair comparison, we applied a similar setting for CLIP-FT, and report the results in Table 3. Our proposed DPOD consistently outperforms the state-of-the-art CLIP-FT for all the unseen domain categories.
Handling Domain-Label Inconsistencies. Often, there are inconsistencies in the domain labels, which may be due to crowd-sourcing for obtaining labels. For example, in the NewsCLIPpings dataset, there are domain labels like football, sports, sport and also Healthcare Network, Healthcare Medicine, etc. which should ideally have the same domain labels. Fig. 1(b) shows the cosine-similarity matrix that is calculated by taking the cosine similarity between the learnt domain-specific prompts of the corresponding two domains. We observe that the learnt domain-specific prompts inherently captures the semantic information of the domains, thereby bringing the similar domains closer, thus helping in their training. Our proposed DPOD does not require the exact domain names, and thus can work even if it is difficult to specify the domains, and dummy domain labels (like $D_{1},D_{2},etc.$) are used instead.
Ablation Studies. Here, we analyze the importance of Stage 1 (label-aware alignment) and Stage 2 (semantic domain vectors) for the all domain scenario in Table 5. We observe that removing either of the stages deteriorates the overall performance. The second row in this table is generated by initializing $\{{V}_{1},{V}_{2},{V}_{3}\}$ with “A photo of” and appending it with a fourth vector ${V}_{4}$, which is obtained by passing a unique one-hot vector corresponding to each training domains. This also implies that if the domains are treated independently, then they cannot fully benefit from the training examples of closely related domains as in the proposed DPOD. The importance of domain-specific prompts can also be observed from Table 5, which shows the performance of prompt $\{{V}_{1},{V}_{2},{V}_{3}\}$ and $\{{V}_{1},{V}_{2},{V}_{3},{V}_{4}\}$, where ${V}_{4}$ is another learnable prompt, but not a domain specific one as in DPOD (third row). Thus, all the proposed modules contribute to the improved performance of DPOD framework.
Table 4: Usefulness of the two stages of the DPOD framework. Effectiveness of Stage 2 was evaluated by taking unique one-hot vectors corresponding to the unique domains.
Table 5: Usefulness of domain-specific prompt tuning. Here the models have been trained on data from all the domains.
@@
$\times$ 68.205
$\times$ 68.923
@@
Politics Sport All Domain
{${V}_{1}$, ${V}_{2}$, ${V}_{3}$} 77.027 69.68 69.349
{${V}_{1}$, ${V}_{2}$, ${V}_{3}$, ${V}_{4}$} 75 68.617 70.351
@@
Table 5: Usefulness of domain-specific prompt tuning. Here the models have been trained on data from all the domains.
Qualitative Results and Limitations. Here, we provide some qualitative results of our framework. Fig. 3 (left with green border) shows few examples of correct predictions. We observe that the model demonstrates the capability to accurately assess the veracity of a news item even in complex situations, as in the first example of a promotional event for a different brand. Fig. 3 (right with pink border) shows few examples of incorrect predictions by our model. We observe that in some cases, the model struggles when multiple elements mentioned in the caption are present in the image, as Koala in the second failure case. In Fig. 4, we present a few qualitative results illustrating where A-CLIP focuses on in both the image and text after Stage 1 in the proposed DPOD method. We observe that, in most cases, A-CLIP attends to regions (marked in yellow) that play a crucial role in decision making.

### v4

@@ 4.2 Stage 2: Computing Semantic Domain Vectors
Here, we create semantic domain vectors which represent how similar each domain is to all the domains in the training dataset. If the number of domains in the training dataset $\mathcal{S}_{train}$ is $n$, we create $n$ semantic domain vectors each having $n$ indices. Let us take the “politics” domain in the training dataset as an example. To compute the semantic domain vector for “politics”, we pass all its images and corresponding text captions, i.e., $\{({I}_{k},{T}_{k})\},k=1,2,...,{n}_{{D}_{i}}$ through the frozen encoders of the A-CLIP trained in Stage 1 to get $\{({\hat{I}}_{k},{\hat{T}}_{k})\}$. Here, ${n}_{{D}_{i}}$ denotes the number of training samples in the politics domain. We take a Hadamard product of the image and text embedding $\{({\hat{I}}_{k},{\hat{T}}_{k})\}$ to get ${n}_{{D}_{i}}$ joint embedding. Then, we compute the mean across all the sample joint embeddings to get the final mean joint embedding $\mathcal{J}_{{D}_{i}}$, particular to a domain ${D}_{i}$ (here, “politics”) as shown below:
@@
This is illustrated in Fig. 2. Finally, to create the semantic domain vector ${w}_{{D}_{i}}$, we take the cosine similarity between this joint embedding $\mathcal{J}_{{D}_{i}}$ and those of the other training domains $\mathcal{J}_{{D}_{k}},k=1,2,...,n$. Hence, the $j^{th}$ index of the ${w}_{{D}_{i}}$ vector will be the cosine similarity of $\mathcal{J}_{{D}_{i}}$ with $\mathcal{J}_{{D}_{j}}$, ($j\in k$) as shown in eq. (6).
@@ 4.3 Stage 3: Domain-Specific Prompt-tuning with Out-of-Domain data
Next, we utilize the semantic domain vectors to learn domain-specific prompt embedding that help to incorporate domain-specific information in the model while classifying a news item as true or fake. We initialize a prompt prefix with the standard text “A photo of” [20]. However, we make this prefix learnable (similar to [31]), since it may not be the best choice for our application. We denote these generic learnable prompts as $\{{V}_{1},{V}_{2},{V}_{3}\}$. To incorporate the domain information, we pass the corresponding semantic domain vector ${w}_{{D}_{i}}$ (obtained in the previous stage) through a linear layer, denoted by $\mathcal{F}_{\theta}$. This layer $\mathcal{F}_{\theta}$ projects the ${n}_{{D}_{i}}$-dimensional domain vector to a 512-dimensional domain prompt token $\mathcal{V}_{{D}_{i}}$, which is then appended to the prompts $\{{V}_{1},{V}_{2},{V}_{3}\}$. This forms the whole learnable prompt structure, where $\{{V}_{1},{V}_{2},{V}_{3}\}$ is trained in a domain-agnostic manner, whereas, $\mathcal{V}_{{D}_{i}}$ is trained by conditioning on the domain of the particular training sample. This entire prompt is then prepended to the text caption ${T}$ and passed through the frozen A-CLIP text encoder to get a text embedding $\mathcal{\hat{T}}$ as shown in eq. (7).
@@
The Classifier Network consists of two fully connected layers which projects $\mathcal{\hat{J}}$ to $\mathcal{\hat{X}}$. Inspired by [7], we also introduce a residual connection from $\mathcal{\hat{J}}$ to $\mathcal{\hat{X}}$, the output of which finally passes through two fully connected layers to give a final prediction $\hat{y}$. The whole network is trained in an end-to-end manner using a Binary Cross-Entropy loss as shown in eq. (8).
@@
Difference with other prompt-learning techniques. As discussed in the Related Work section, there are some existing works on domain-based prompt tuning [8, 23]. But there are few important differences between these techniques and the proposed one:
1) In the existing works, “domain” or “style” refers to changes in image features. Contrary to this, in our case, domains (such as, “politics”, “law-crime”, etc.) do not have a significant change in image features (since, images from multiple domains will contain similar objects like humans, etc.), but are subjective to human interpretation of both the images and the texts jointly.
@@
Table 1: Performance comparison of the proposed DPOD framework with existing approaches on the NewsCLIPpings dataset, where we train and test on all the domains.
Methods Backbone Accuracy (%) DT-Transformer [17] CLIP ViT-B/32 65.67 MNSL [28] CLIP 68.20 DCD [10] CLIP 65.40 NewsCLIPpings [15] CLIP ViT-B/32 60.23 CLIP-RN50 61.62 CLIP-FT [1] CLIP ViT-B/16 64.79 CLIP ViT-B/32 65.10 SSDL [16] CLIP ViT-B/16 65.00 CLIP RN-50 71.00 DPOD (Ours) CLIP ViT-B/16 70.49 DPOD (Ours) CLIP ViT-B/32 70.81 CLIP RN-50 74.43
@@ 5 Experimental Evaluation
First, we evaluate the effectiveness of the proposed framework on the standard large-scale benchmark, namely NewsCLIPpings dataset [15], which is created from VisualNews dataset [14]. It contains news articles from sources like: The Guardian, BBC, USA Today, and The Washington Post, split into $71,072$ train, $7,024$ val and $7,264$ test examples. It contains news articles from 54 domains with varied amounts of training data across each of the domains, e.g., politics (1396 samples), law-crime (4348 samples), sports (2786 samples) etc., reflecting the imbalance problem.
Implementation Details. We used CLIP ViT-B/32[20] model as the backbone for all our experiments. We use accuracy as the evaluation metric as in [1, 16], as the data has equal distribution of true and fake news. Our codes are implemented in PyTorch and uses a single NVIDIA GeForce RTX 3080 Ti GPU. ADAM optimizer with weight decay and stability constant is used for the training. All the results are reported using the same set of parameters, e.g., $\beta=1.5$, $lr=1e-4$ and $batchsize=64$ across all dataset splits. The values of hyperparameters are obtained during the validation process against the constant validation dataset containing all-domain data. These values were then kept fixed for the rest of the experiments. Stage 1 is run for 40 epochs, which remains fixed even if the desired domain changes. We initialize our domain-agnostic prompts $\{V_{1},V_{2},V_{3}\}$ required in Stage 3, as “A photo of” which is then learnt through training. After obtaining the output probability at the end of Stage 3, we use Sigmoid function to convert the output into a probability score $\hat{y}$ between 0 to 1. We train Stage 3 for 30 epochs. At the time of inference, the data having output greater than or equal to $0.5$ is assigned the label 1, i.e., Fake and the others are assigned label 0, i.e., True. Now, we describe the experiments conducted to evaluate the effectiveness and generalization of the proposed DPOD framework. We would specifically like to answer the following research questions:
i) How does the model perform in the standard setting, where all the domain data is used for training and testing?
ii) How well does the proposed model perform for a desired target domain?
iii) Can our model generalize to completely unseen target domains and to other datasets?
iv) Can the model handle inconsistencies in domain labels and how well does it understand about the domains?
v) Are all the components in the model useful?
@@
Radio
Education
Money
Travel
@@
Evaluation on standard setting for all domains: DPOD is a generalized framework, where the model is trained with data from all the available domains. First, we evaluate DPOD on the standard test protocol, where image-text pairs are available from all the domains, and report the results in Table 1. Here, we assume that we know the domain label of the test data. The experiment is run $5$ times with random sampling of training data, and the mean performance is reported. We compare our proposed framework with the previous SOTA approaches on NewsCLIPpings dataset, namely, CLIP-FT [1], SSDL [16], DT-Transformer [17], MNSL[28], DCD[10], NewsCLIPpings[15]. For CLIP-FT, the entire CLIP model is fine-tuned along with two additional layers with the available training data. This is the performance without using the open-domain data. For all the approaches, we directly report the accuracies from the respective papers. When domain knowledge of the test sample is unknown, DPOD gives an accuracy of $69.13\%$, which is comparable to the performance achieved when the test domain information is known. For fair comparison, we also perform our experiments with different variants of CLIP, namely, ViT-B/32, ViT-B/16 and RN-50. We observe that DPOD consistently outperforms the SOTA by a significant margin when similar backbones are used for comparison.
@@
Figure 3: Similarity of the learnt domain-specific prompts of different domains. We observe that the sets Football, Sport, Sports and Healthcare Network, Healthcare Medicine have similar domain-specific prompts among themselves, but they have low similarity with each other and also with Music.
@@
Evaluation on a specific domain: Now, we consider the scenario where an individual is interested in news from a specific domain, eg a doctor might be interested in healthcare news. Unlike existing methods, DPOD simultaneously specializes for each domain, by effectively utilizing out-of-domain samples. To evaluate how the model handles data imbalance (from different domains) in the training data, we report results on $5$ domains with most datapoints and $5$ domains with the least data samples in Table 2. In the absence of comparative approaches, we experiment with the following strong baselines: i) Clip-FT [1]: without the open-domain information for fair comparison with the other approaches. ii) CoOp [31]: First, the prompts $\{V_{1},V_{2},V_{3}\}$ are learnt using training data from all the domains. Then, these prompts are frozen for evaluation on the target domains. For fair comparison with our approach, CoOp is also used with A-CLIP. All the models here are trained on the entire dataset, and then evaluated on the domain of interest. We observe that DPOD consistently outperforms the other methods for all the domains, irrespective of the number of training examples available for that domain.
Generalization to unseen domains. In this dynamic world, the model may encounter test data, which it has not been trained on. Here, we evaluate the performance of our model when presented with completely new domains not included in the training set. For this, we train our model using samples from 90% of the domains, and test it on the remaining 10% unseen domains, using the most similar domain-specific prompt (based on cosine similarity). For a fair comparison, we applied a similar setting for CLIP-FT, and report the results in Fig. 4. Here, the most similar domains corresponding to the 5 unseen domains obtained are as follows: politicsintl relns(sim=0.61), law crimeconflict attack(sim=0.78), arts cult.art & design(sim=0.59), musicmedia(sim=0.76), sportsfootball(sim=0.66). Our proposed DPOD consistently outperforms the state-of-the-art CLIP-FT for all the unseen domain categories.
@@
Figure 4: Comparison of accuracies (%) where we train the model on 90% domains and evaluate on the rest 10% unseen domains.
@@
$\times$ $\checkmark$ $68.205$
$\checkmark$ $\times$ $68.923$
@@
Table 3: Usefulness of the two stages of the proposed DPOD framework. Effectiveness of Stage 2 was evaluated by passing unique one-hot vectors corresponding to the unique domains.
@@
Politics Sport All Domains
{${V}_{1}$, ${V}_{2}$, ${V}_{3}$} $77.027$ $69.680$ $69.349$
{${V}_{1}$, ${V}_{2}$, ${V}_{3}$, ${V}_{4}$} $75.000$ $68.617$ $70.351$
@@
Table 4: Usefulness of domain-specific prompt-tuning. Here the models have been trained on data from all the domains.
Handling Domain-Label Inconsistencies. Often, there are inconsistencies in the domain labels, which may be due to crowd-sourcing for obtaining labels. For example, in the NewsCLIPpings dataset, there are domain labels like football, sports, sport and also Healthcare Network, Healthcare Medicine, etc. which should ideally have the same domain labels. The matrix in Fig. 3 shows the cosine similarity between the learnt domain-specific prompts of the corresponding two domains. We observe that the learnt domain-specific prompts inherently captures the semantic information of the domains, thereby bringing the similar domains closer, thus helping in their training. The proposed DPOD does not require the exact domain names, and thus can work even if it is difficult to specify the domains, and dummy domain labels (like $D_{1},D_{2},etc.$) are used instead.
Cross Dataset Performance. To further test the generalization capability of DPOD, we train our model and evaluate it on another real-world dataset, VERITE [18], which lacks domain information. For this, we create 10 clusters from the joint image-text embeddings using k-means, and create domain similarity vectors corresponding to these 10 cluster means. We then pass this through our trained model to create domain-specific vectors. We report the performances in Table 5. We observe that DPOD performs significantly better than the other methods in a similar setting.
Method Backbone Accuracy (%)
DT-Transformer [17] CLIP ViT-L/14 $57.50$
CLIP-FT [1] CLIP ViT-B/32 $56.78$
DPOD CLIP ViT-B/32 62.15
Table 5: Cross dataset performance. All methods are trained on NewsCLIPpings and evaluated on VERITE. DPOD outperforms other methods including DT-Transformer of a superior backbone.
Ablation Studies. Here, we analyze the importance of Stage 1 (label-aware alignment) and Stage 2 (semantic domain vectors) for the all domains scenario in Table 3. We observe that removing either of the stages deteriorates the overall performance. The second row in this table is generated by initializing $\{{V}_{1},{V}_{2},{V}_{3}\}$ with “A photo of” and appending it with a fourth vector ${V}_{4}$, which is obtained by passing a unique one-hot vector corresponding to each training domains. This implies that if the domains are treated independently, they cannot fully benefit from the training examples of closely related domains as in DPOD. The importance of domain-specific prompts can also be observed from Table 4, which shows the performance of prompt $\{{V}_{1},{V}_{2},{V}_{3}\}$ and $\{{V}_{1},{V}_{2},{V}_{3},{V}_{4}\}$, where ${V}_{4}$ is another learnable prompt, but not a domain specific one as in DPOD (third row). Thus, all the proposed modules contribute to the improved performance of DPOD framework.
Qualitative Results and Limitations. Here, we provide some qualitative results. Fig. 5 (left with green border) shows few examples of correct predictions. We observe that the model shows the ability to accurately assess the veracity of a news item even in complex situations, as in the first example of a promotional event for a different brand. Fig. 5 (right with pink border) shows few examples of incorrect predictions by our model. We observe that in some cases, the model struggles when multiple elements mentioned in the caption are present in the image, as Koala in the second failure case. In Fig. 6, we show some qualitative results of the A-CLIP attention maps on both the image and the text. We observe that, generally A-CLIP attends to regions (in yellow) that play a crucial role in decision making.